OpenStudy (anonymous):

Find all the values of k so that the trinomial can be factored using integers. x^2+kx-19

OpenStudy (anonymous):

Use the discriminant \[\sqrt{b^2-4ac}\]

OpenStudy (anonymous):

whats that....

OpenStudy (anonymous):

It says integers, so it has to be real.

OpenStudy (anonymous):

well it says integers, so they have to be integers

OpenStudy (anonymous):

The discriminant has to equal to zero

OpenStudy (anonymous):

So \[k^2+4(19)=0\]

OpenStudy (anonymous):

there are not too many factors of \(-19\)

OpenStudy (anonymous):

The discriminant is without the square root sign sorry.

OpenStudy (anonymous):

it is either \((x+19)(x-1)\) or \((x-19)(x+1)\)

OpenStudy (anonymous):

so the thing would be 19,-19?

OpenStudy (anonymous):

multiply these two out, see what you get for the middle term

OpenStudy (anonymous):

lol i dont get it

OpenStudy (anonymous):

you have \(x^2+kx-19\) and you want to know what \(k\) is so that this will factor so it has to look like either \((x+19)(x-1)\) which is one way to factor, or else it must look like \((x-19)(x+1)\) there is no other way because there are not any more integers whose product is \(-19\)

OpenStudy (anonymous):

multiply the first one out and you get \[(x+19)(x-1)=x^2+19x-x-19=x^2+18x-19\] so \(k\) could be \(18\)

OpenStudy (anonymous):

your last job is to multiply out \[(x-19)(x+1)\]and see what you get for the middle term